# Prompt Sensitivity and Answer Consistency of Small Open-Source Large Language Models on Clinical Question Answering: Implications for Low-Resource Healthcare Deployment

### 04 Mar 2026 v2

@@ Abstract
Methods: Five open-source language models were evaluated across three clinical question answering datasets (MedQA, MedMCQA, PubMedQA) using five prompt styles. Consistency scores, accuracy, and instruction-following failure rates were measured using experiments conducted locally on consumer CPU hardware without fine-tuning.
Results: Consistency and accuracy were largely independent across models and datasets. Gemma 2 achieved the highest consistency scores (0.845–0.888) but the lowest accuracy (33.0–43.5%), while Llama 3.2 demonstrated moderate consistency (0.774–0.807) alongside the highest accuracy (49.0–65.0%). Roleplay prompts consistently reduced accuracy across all models and datasets, with Phi-3 Mini showing the largest decline of 21.5 percentage points on MedQA. Instruction-following failure rates varied by model and were not determined by parameter count, with Phi-3 Mini exhibiting the highest UNKNOWN rate at 10.5% on MedQA. Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure on PubMedQA (99.0% UNKNOWN rate), demonstrating that domain knowledge alone is insufficient for structured clinical question answering.
Conclusions: High consistency does not imply correctness in small clinical language models; models can be reliably incorrect, representing a particularly dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among the models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
@@ 1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, from clinical decision support to exam-style medical question answering. While much attention has focused on large, cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). However, as these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but also on how stable and predictable their outputs are under different conditions. Despite rapid adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
@@
Most existing evaluations of clinical language models focus primarily on accuracy, that is, whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction matters. A model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
@@
To our knowledge, no prior study has systematically evaluated whether small, CPU-runnable language models provide consistent answers under controlled prompt variation in clinical settings. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Recent surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs without analyzing answer consistency as a reliability dimension. The present study addresses this gap directly.
@@
The present work systematically evaluates prompt sensitivity in five open-source language models, including Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained), across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA) using five controlled prompt styles designed to simulate realistic query reformulation. A quantitative consistency score is introduced that measures how often a model produces the same answer across prompt variations for the same question and analyzes this alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare settings.
@@
Four central findings emerge from the analysis. First, consistency and accuracy are independent metrics: models that produce highly stable answers across prompt variations are not necessarily more correct, with Gemma 2 demonstrating the clearest case of high consistency alongside low accuracy. Second, roleplay prompt styles systematically reduce performance across all models and datasets, indicating that persona-based framing negatively affects clinical question answering reliability. Third, instruction-following failure rates vary independently of model size, demonstrating that larger parameter counts do not guarantee safer or more dependable outputs under constrained deployment conditions. Fourth, Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction- following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical deployment.
@@ 3.1 Datasets
From each dataset, 200 questions were randomly sampled using a fixed random seed (seed = 42) to ensure reproducibility, yielding 600 questions total across three distinct clinical reasoning contexts.
@@ 3.2 Models
Five open-source language models spanning a parameter range of 2B to 7B parameters were evaluated, including one domain-adapted medical pretraining model to examine the effect of model scale on prompt consistency. The selected models were Phi-3 Mini (Abdin and others, 2024) (3.8B parameters, Microsoft), Llama 3.2 (Grattafiori and others, 2024) (3B parameters, Meta), Gemma 2 (Gemma Team and others, 2024) (2B parameters, Google), and Mistral (Jiang and others, 2023) (7B parameters, Mistral AI). These models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families, enabling cross-architecture comparison of consistency behavior. All inference was conducted locally using Ollama without any domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behavior rather than task-specific adaptation.
@@
In addition to these instruction-tuned general models, Meditron-7B (Chen and others, 2023), a domain-adapted language model developed by EPFL, was included. Meditron-7B is built on Llama-2 and further pretrained on a curated medical corpus comprising PubMed articles, biomedical abstracts, and internationally recognized clinical practice guidelines. Critically, Meditron-7B is a base pretraining model without instruction tuning, meaning it has not been trained to follow structured task instructions. It was explicitly included to examine whether domain-specific medical pretraining alone, in the absence of instruction tuning, is sufficient for structured clinical question answering and to contrast its behavior with that of general instruction-tuned models. This comparison addresses a practical question for clinical AI deployment: whether domain knowledge alone is sufficient or whether instruction-following capability is equally essential.
@@ 3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures fully deterministic outputs, meaning the model produces identical responses for identical prompts. This is essential for isolating prompt sensitivity as the source of answer variation rather than stochastic sampling. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures. All responses, including UNKNOWN outputs, were retained for consistency scoring and analysis.
@@ 3.5 Evaluation Metrics
Per-Style Accuracy. Accuracy is computed separately for each prompt style (original, formal, simplified, roleplay, direct), enabling identification of which formulations the model handles most and least reliably.
@@
Statistical significance was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. All p-values are reported unadjusted. Given the exploratory nature of this analysis, no correction for multiple comparisons was applied.
Together, these metrics provide a holistic assessment of model behavior. A model may be highly consistent but consistently wrong, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
@@ 5.1 Consistency and Accuracy are Independent Metrics
A central finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), the results presented here demonstrate that a model can exhibit high consistency while still producing systematically incorrect answers, a failure mode referred to in this work as reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically incorrect.
@@ 5.5 Limitations
Several limitations should be considered when interpreting these findings. First, 200 questions per dataset were sampled, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used, MedQA, MedMCQA, and PubMedQA, focus primarily on structured examination-style and research-based questions and may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
@@
Together, these limitations suggest that the results should be interpreted as an initial systematic analysis of prompt sensitivity under constrained deployment conditions rather than a comprehensive evaluation of all clinical AI scenarios. Future work should address these gaps through larger samples, broader dataset coverage, and human clinical validation studies.

### 16 Mar 2026 v3

# Prompt Sensitivity and Answer Consistency in Small Open-Source Language Models for Clinical Question Answering: An Empirical Evaluation for Low-Resource Healthcare Deployment

@@ Abstract
Methods: Five open-source language models spanning distinct architectural and training paradigms were evaluated across three clinical question answering datasets (MedQA, MedMCQA, PubMedQA) using five controlled prompt style variations, yielding 15,000 total inference calls conducted locally on consumer CPU hardware without fine-tuning. Consistency scores, accuracy, and instruction-following failure rates were measured and interpreted in the context of each model’s architectural design.
Results: Consistency and accuracy were largely independent across models and datasets. Gemma 2 achieved the highest consistency scores (0.845–0.888) but the lowest accuracy (33.0–43.5%), producing perfectly consistent yet incorrect answers on 77 of 200 MedQA questions (38.5%), a failure mode termed reliable incorrectness. Llama 3.2 demonstrated moderate consistency (0.774–0.807) alongside the highest accuracy (49.0–65.0%). Roleplay prompts consistently reduced accuracy across all models and datasets, with Phi-3 Mini showing the largest decline of 21.5 percentage points on MedQA. Instruction-following failure rates varied by model and were not determined by parameter count, with Phi-3 Mini exhibiting the highest UNKNOWN rate at 10.5% on MedQA. Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure on PubMedQA (99.0% UNKNOWN rate), demonstrating that domain knowledge alone is insufficient for structured clinical question answering.
Conclusions: High consistency does not imply correctness in small clinical language models; models can be reliably incorrect, representing a potentially dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among the models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
@@ 1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, including clinical decision support, medical documentation, and exam-style clinical question answering. While much attention has focused on large cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). As these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but also on how stable and predictable their outputs are under different conditions. Despite increasing adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
@@
Most existing evaluations of clinical language models focus primarily on accuracy, that is, whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction is important. A model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
@@
Prior work has not systematically evaluated whether small, CPU-runnable language models provide consistent answers under controlled prompt variation in clinical contexts. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs but do not analyze answer consistency as a reliability dimension. The present study addresses this gap.
This study addresses three research questions:
RQ1: Do small open-source language models produce consistent answers when semantically equivalent clinical questions are phrased differently?
RQ2: Is answer consistency correlated with clinical question answering accuracy?
RQ3: How do different prompt styles influence the reliability of model outputs?
@@
To investigate these questions, this work evaluates five open-source language models—Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained)— across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA). Each question is evaluated under five controlled prompt styles designed to simulate realistic query reformulation. A quantitative consistency score is introduced to measure how often a model produces the same answer across prompt variations for the same question, and this metric is analyzed alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare environments.
@@
The analysis yields several key findings. First, consistency and accuracy emerge as independent performance dimensions: models that produce highly stable answers across prompt variations are not necessarily more accurate. Second, roleplay prompt styles consistently reduce performance across models and datasets, indicating that persona-based prompting may degrade reliability in clinical question answering tasks. Third, instruction-following failures vary independently of model size, suggesting that larger parameter counts do not guarantee more dependable outputs. Finally, Meditron-7B, a domain-pretrained model without instruction tuning, exhibits near-complete instruction-following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are distinct requirements for clinical AI deployment.
@@ 3 Methods
This study follows a controlled empirical evaluation design to assess the reliability of small language models under prompt variation in clinical question answering tasks.
@@ 3.1 Datasets
From each dataset, 200 questions were randomly sampled using a fixed random seed (seed = 42) to ensure reproducibility. This sample size was selected to balance statistical coverage with the computational constraints of local CPU inference across multiple models and prompt variations, resulting in 15,000 total inference calls.
@@ 3.2 Models
Five open-source language models spanning 2B–7B parameters were evaluated, including one domain-adapted medical pretraining model. Table 1 summarises the architectural and training characteristics of each model. The four instruction-tuned models differ in their fine-tuning methodology, pretraining scale, and prompt template conventions in ways that plausibly influence prompt sensitivity behaviour.
Phi-3 Mini (Abdin and others, 2024) (3.8B, Microsoft) is a dense decoder-only transformer fine-tuned using Supervised Fine-Tuning (SFT), Direct Preference Optimisation (DPO), and Reinforcement Learning from Human Feedback (RLHF). Its pretraining relies heavily on high-quality synthetic data and filtered educational content, an approach motivated by the “textbooks are all you need” philosophy. This regime prioritises structured reasoning and instruction adherence; however, the model’s strict chat template format (<|user|>...<|end|>) may create sensitivity to prompts that deviate from this expected structure, contributing to the elevated UNKNOWN rate observed on MedQA.
Llama 3.2 (Grattafiori and others, 2024) (3B, Meta) is derived from Llama 3.1 via pruning and knowledge distillation from larger Llama 3.1 models (8B and 70B), followed by multiple rounds of post-training alignment using SFT, Rejection Sampling (RS), and DPO. This multi-round alignment process, explicitly optimising dialogue helpfulness across varied instruction formats, likely contributes to the model’s strong instruction-following performance and its low UNKNOWN rates across all datasets.
Gemma 2 (Gemma Team and others, 2024) (2B, Google DeepMind) is trained via knowledge distillation from a substantially larger teacher model rather than standard next-token prediction, using on-policy distillation to mitigate train-inference distribution mismatch. Post-training applies SFT, RLHF with a conversational reward model, and model merging across hyperparameter configurations. Knowledge distillation at this scale may instil stable but potentially miscalibrated response tendencies inherited from the teacher model, providing a mechanistic account of Gemma 2’s high consistency scores alongside its low accuracy: the model may be replicating confident response patterns from its teacher that do not generalise correctly to clinical reasoning tasks.
Mistral (Jiang and others, 2023) (7B, Mistral AI) uses sliding window attention and grouped-query attention for efficient inference, and is instruction-tuned using SFT on publicly available conversation datasets. Compared to Llama 3.2 and Phi-3, Mistral’s instruct fine-tuning used a more limited set of public instruction data and requires strict [INST]...[/INST] delimiters. Prompts that do not conform to this format may partially escape instruction alignment, potentially contributing to higher UNKNOWN rates relative to Llama 3.2 and Gemma 2.
@@
Meditron-7B (Chen and others, 2023) (EPFL) is built on Llama-2 and further pretrained on a curated medical corpus comprising PubMed abstracts and international clinical practice guidelines. Critically, Meditron-7B is a continued-pretraining model without instruction tuning; it has not been trained to follow structured task instructions or produce constrained single-token outputs. Its inclusion serves as an architectural control condition: by contrasting a domain-knowledgeable but non-instruction-tuned model against instruction-tuned general models, this study isolates instruction-following capability as a distinct and necessary architectural requirement for structured clinical question answering, independent of domain knowledge.
All models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families with meaningfully different training methodologies, enabling cross-architecture comparison of consistency behaviour. All inference was conducted locally using Ollama without domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behaviour rather than task-specific adaptation.
[…]
Model Params Base Model Inst. Tuned Fine-Tuning Method Medical Domain Chat Template
Phi-3 Mini 3.8B Phi-3 (Microsoft) Yes SFT + DPO + RLHF No <|user|>..<|end|>
Llama 3.2 3B Llama 3.1 (Meta) Yes SFT + RS + DPO (multi-round) No <|start_header_id|>..<|eot_id|>
Gemma 2 2B Gemma 2 (Google) Yes KD + SFT + RLHF + Model Merging No <start_of_turn>..<end_of_turn>
Mistral 7B 7B Mistral (Mistral AI) Yes SFT (public datasets) No [INST]..[/INST]
Meditron-$\dagger 7B$ Llama-2 (Meta) No Continued pretraining only (PubMed + guidelines) Yes None
@@ 3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures deterministic outputs, meaning that identical prompts produce identical responses across runs. This configuration isolates prompt variation as the primary source of output differences rather than stochastic sampling effects. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures where the model does not produce a valid answer option. All responses, including UNKNOWN outputs, were retained for consistency scoring and subsequent analysis. All inference parameters were held constant across models and prompt styles to ensure fair comparison.
@@ 3.5 Evaluation Metrics
Per-Style Accuracy. Accuracy is computed separately for each prompt style (original, formal, simplified, roleplay, direct), enabling identification of which prompt formulations the model handles most and least reliably.
@@
Statistical significance was assessed using the Wilcoxon signed-rank test for comparisons of consistency scores and McNemar’s test for paired accuracy comparisons. All p-values are reported without adjustment for multiple comparisons. Because the analysis is exploratory and hypothesis-generating, no correction for multiple testing was applied, consistent with prior exploratory clinical AI benchmarking studies (Bedi and others, 2025). Individual p-values should therefore be interpreted cautiously rather than as confirmatory statistical evidence.
Together, these metrics provide a holistic assessment of model behavior. A model may be highly consistent yet systematically incorrect, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
@@ 4.3 Instruction-Following Failure Rate
The variation in UNKNOWN rates across instruction-tuned models also reflects differences in the scope and methodology of their post-training alignment. Llama 3.2 and Gemma 2, which produced the lowest UNKNOWN rates, underwent multi-round alignment processes optimising across diverse instruction formats. Mistral 7B and Phi-3 Mini, which showed higher UNKNOWN rates on certain datasets, relied on more constrained instruction fine-tuning that may be more sensitive to prompt phrasings that diverge from their expected template structures. These patterns suggest that the breadth of instruction format coverage during fine-tuning, not merely parameter count, is the primary architectural determinant of instruction-following reliability in structured clinical tasks.
@@ 5 Discussion
From a clinical informatics perspective, these findings highlight the importance of evaluating AI systems not only for accuracy but also for response stability. Clinical decision support tools are often queried in multiple ways by clinicians, and inconsistent responses to semantically equivalent queries may undermine trust in AI-assisted systems. Conversely, consistently incorrect outputs may create a false sense of reliability if repeated answers are interpreted as evidence of model confidence. Evaluation frameworks that jointly assess accuracy and response stability therefore provide a more comprehensive basis for assessing whether language models are sufficiently reliable for integration into clinical workflows.
@@ 5.1 Consistency and Accuracy are Independent Metrics
A central empirical finding of this study is that consistency and accuracy represent independent dimensions of small language model performance in clinical question answering. While prior evaluations of clinical AI have predominantly focused on accuracy as the primary metric (Bedi and others, 2025), the results presented here demonstrate that a model can exhibit high consistency while still producing systematically incorrect answers, a failure mode referred to in this work as reliable incorrectness.
Gemma 2 illustrates this most clearly, achieving the highest mean consistency scores across all datasets (0.845–0.888) while recording the lowest overall accuracy (33.0–43.5%). This pattern suggests that the model has learned stable response tendencies that are nonetheless clinically incorrect. Across 200 MedQA questions, Gemma 2 produced perfectly consistent yet incorrect answers on 77 questions (38.5%), selecting the same wrong option across all five prompt styles with a consistency score of 1.0.
@@
The architectural characteristics of Gemma 2 offer a plausible mechanistic account of this pattern. The 2B Gemma 2 model is trained via knowledge distillation from a substantially larger teacher model, inheriting the teacher’s response tendencies rather than learning independently from ground-truth labels. If the teacher model has developed systematic biases in clinical question answering — for instance, preferring certain option positions or applying heuristics that generalise poorly to medical reasoning — these biases may be faithfully replicated in the distilled student model, producing high cross-prompt consistency that reflects the teacher’s error patterns rather than genuine clinical knowledge. This distillation-induced miscalibration represents a distinct architectural risk factor for clinical AI deployment that accuracy-only evaluation frameworks would not detect.
@@ 5.2 Roleplay Prompts Consistently Underperform
This phenomenon is likely amplified by the prompt template conventions enforced during instruction fine-tuning. Each model evaluated uses a distinct chat template: Phi-3 Mini employs <|user|>...<|end|> delimiters, Llama 3.2 uses <|start_header_id|>...<|eot_id|> markers, and Mistral 7B requires [INST]...[/INST] framing. Roleplay prompts, which begin with persona instructions such as “You are a senior physician,” may partially misalign with the expected input structure of these templates, shifting the model’s processing toward narrative or conversational generation rather than constrained structured reasoning. This suggests that roleplay sensitivity reflects not merely task framing but an interaction between prompt style and the learned template expectations embedded during instruction tuning — an architectural rather than purely behavioural phenomenon.
@@ 5.4 Domain Knowledge Versus Instruction Following
From an architectural standpoint, this finding reflects a fundamental distinction between two training objectives that are often conflated in clinical AI discourse. Continued pretraining on medical corpora, as applied to Meditron-7B, optimises the model’s internal representations of medical knowledge but does not instil the structured output behaviour required for task-constrained deployment. Instruction tuning, by contrast, explicitly trains the model to produce constrained responses in specific formats in response to specific prompt structures. These are orthogonal capabilities: a model can encode rich biomedical knowledge through pretraining yet remain clinically unusable because it lacks the architectural conditioning to translate that knowledge into a structured, deployable output. For practitioners selecting models for low-resource clinical deployment, this distinction implies that medical domain pretraining alone is an insufficient selection criterion; instruction-following capability must be verified independently.
@@ 5.5 Limitations
Several limitations should be considered when interpreting these findings. These limitations define the scope of the current study and collectively motivate the future directions described in Section 5.6.
First, 200 questions per dataset were sampled, which provides a computationally feasible and reproducible evaluation but may not fully capture the diversity of each benchmark. Second, the three datasets used, MedQA, MedMCQA, and PubMedQA, focus primarily on structured examination-style and research-based questions and may not reflect the complexity of real-world clinical dialogue or unstructured patient interactions.
@@
Together, these limitations define the scope of this study as an initial systematic analysis of prompt sensitivity under constrained deployment conditions. The evaluation framework introduced here is designed to be reproducible and extensible, supporting future clinical validation studies without requiring specialised infrastructure.